The current in an RL circuit increases to 98 % of its final value 2.22 safter the switch is closed.

PART A)

What is the time constant for this circuit?

PART B)

If the inductance in the circuit is 0.298 H , what is the resistance?
Answer #1

I = I0( 1-e - t/\tau )

\tau, time constant = L/R , t = time in second =2.22s

Io is full current

0.98 =1(1-e-2.22/\tau)

e-2.22/\tau = 0.02

taking loge

-2.22/\tau = loge 0.02

\tau = 0.56748 sec part A answere

part B

\tau =L/R

0.56748 = 0.2986/R

R= 0.52618 \Omega